# A Random Rule Model

### 04 Mar 2026 v1

@@ 2.1 Setup
A menu (choice set) is a binary set $A=\{L^{1},L^{2}\}$ of lotteries, each lottery being a finite-support distribution over monetary outcomes. We label $L^{1}$ the left choice and $L^{2}$ the right choice. Let $\mathcal{A}_{\mathrm{full}}$ denote the full set of menus in the dataset. Each $A\in\mathcal{A}_{\mathrm{full}}$ is observed $n(A)\geq 1$ times. Let $x_{m}(A)\in\{L^{1},L^{2}\}$ be the realized choice in repetition $m\in\{1,\dots,n(A)\}$. [^1] Within each menu we label options as “Left” ($L^{1}$) and “Right” ($L^{2}$) using the dataset’s fixed convention. The menu-level left-choice frequency is
$\hat{p}(A)\ :=\ \frac{1}{n(A)}\sum_{m=1}^{n(A)}\mathbf{1}\{x_{m}(A)=L^{1}\}\in[0,1].$
@@
The primitives of our model are lotteries, and all subsequent rules - including regret- and disappointment-type rules - are defined as functions of the lottery representation (or of a canonical state space constructed from the marginal distributions via the product measure over outcome pairs, as described in Section 4). Hence the lottery representation is without loss of generality for the properties we study. We will use the state-space view only when defining several rules, and otherwise work with the simpler lottery notation introduced above.
@@ 3 Identification via Rule Switching
This section clarifies when the gate parameters $\theta=\{(\alpha_{f},\beta_{f})\}_{f\in\mathcal{F}}$ are pinned down by menu-level choice frequencies. The key identifying force is rule switching: because different rules recommend different sides in different menus, the decisive-side pattern varies across the menu support, and this variation pins down the relative rule weights.
From CA-RRM to a collapsed binary logit index (identity, not a new model).
@@
Let $\mathcal{A}_{2}:=\{A\in\mathcal{A}_{\mathrm{full}}:\ \mathcal{I}_{L}(A)\neq\emptyset,\ \mathcal{I}_{R}(A)\neq\emptyset\}$ denote the set of two-sided menus. For $A\in\mathcal{A}_{2}$, define the population log-odds
@@
Let $\vartheta\in\mathbb{R}^{K}$ collect the remaining free parameters, with $K=(|\mathcal{F}|-1)\,(1+\dim(\psi))$.
@@
and write the odds ratio r(A):=p(A)/(1p(A))r(A):=p(A)/(1-p(A)). For any $x\in\mathbb{R}^{d}$ define the positive weight vector $\omega_{f}(x):=\exp(\alpha_{f}+\beta_{f}^{\top}x)$. Cross-multiplying the odds form of the model yields, for each menu $A$ with $\psi(A)=x$,
@@
i.e. $h(A)^{\top}\omega(x)=0$ where $h_{f}(A):=\kappa_{f}^{L}(A)-r(A)\kappa_{f}^{R}(A)$.
Global identification theorem.
Theorem 1 (Global identification under within-feature activity variation).
Consider the CA-RRM with affine indices $s_{f}(A)=\alpha_{f}+\beta_{f}^{\top}\psi(A)$, $f\in\mathcal{F}$. Assume $p(A)\in(0,1)$ for all menus in the support, and impose the normalization $(\alpha_{f_{0}},\beta_{f_{0}})=(0,0)$. Let $d=\dim(\psi)$. Suppose there exist $(d+1)$ feature values $x^{(0)},\ldots,x^{(d)}$ such that:
@@
    Within-feature variation in decisive-side patterns: For each $k\in\{0,\ldots,d\}$, there exist $M_{k}\geq|\mathcal{F}|-1$ menus $A_{k1},\ldots,A_{kM_{k}}$ with $\psi(A_{km})=x^{(k)}$ for all $m$, and the $M_{k}\times|\mathcal{F}|$ matrix
@@
The proof, given in Appendix B, has two layers. The first layer identifies, at a fixed feature value $x$, the vector of relative rule weights $\omega(x)=(\exp(\alpha_{f}+\beta_{f}^{\top}x))_{f\in\mathcal{F}}$. For any menu $A$ with $\psi(A)=x$, the model says that the odds of choosing left equal the ratio of total weight on rules decisive for left to total weight on rules decisive for right. Cross-multiplying turns this odds identity into a linear equation in the unknown weights $\omega(x)$, with coefficients determined by the decisive-side indicators $(\kappa^{L}(A),\kappa^{R}(A))$ and the observed odds $r(A)$. If we see many menus that share the same feature value $x$ but exhibit different decisive-side patterns (because different rules switch sides across those menus), we obtain many independent linear equations for the same unknown vector $\omega(x)$. Condition (G1) requires that these equations be sufficiently rich - rank $|\mathcal{F}|-1$ - so they pin down the relative weights of all rules at $x$ (the remaining one degree of freedom is an overall scale). The CA location normalization then fixes that scale because it anchors one rule’s weight at $\omega_{f_{0}}(x)\equiv 1$, yielding a unique $\omega(x)$ at each identifying feature value. The second layer then recovers the gate parameters. Once the weights are known at $(d+1)$ feature values, each rule $f$ satisfies the linear relation $\log\omega_{f}(x^{(k)})=\alpha_{f}+\beta_{f}^{\top}x^{(k)}$ across $k$. Condition (G2) (affine independence) ensures that the resulting $(d+1)\times(d+1)$ regressor matrix is invertible, so these $(d+1)$ equations uniquely solve for $(\alpha_{f},\beta_{f})$ rule-by-rule. In short: (G1) uses rule switching at fixed $x$ to identify weights, and (G2) uses variation in $x$ to map those weights back to the affine indices globally, rather than only locally.
Effective dimension.
When the feature vector has structural redundancies (e.g., exact linear dependencies among coordinates), the theorem applies with $d$ replaced by the effective dimension $d_{\mathrm{eff}}:=\mathrm{rank}([\mathbf{1},z])-1$, and identification holds for the gate parameters projected onto the effective feature subspace (i.e., up to the null space of $\mathbf{1},z]$). In practice, one can reparameterize the gate in terms of a $d_{\mathrm{eff}}$-dimensional basis of the column space of $\mathbf{1},z]$, and identification then holds for the parameters associated with this effective feature space. In our application, the baseline gate features $z(A)\in\mathbb{R}^{12}$ have $d_{\mathrm{eff}}=11$ (the expected-value gap is a linear combination of the two level expected values).
Menu encoding.
The feature vector $\psi(A)\in\mathbb{R}^{4M}$ is the fixed-length concatenation of sorted outcomes and probabilities for the two lotteries (Section 6.3 gives the precise definition).
@@
Theorem 1 reduces identification to two verifiable conditions on the observed menu support. We report three diagnostics.
@@
Report the fraction of menus that are two-sided: $\widehat{\Pr}(A\in\mathcal{A}_{2})$. This ensures the log-odds representation (9) is empirically meaningful.
@@
The diagnostic is implemented on the interpretable gate features $z(A)\in\mathbb{R}^{12}$ used in the baseline specification (Section 6.6). To check (G1), group menus into cells that share the same (or nearby) value of $z(A)$, and compute $\mathrm{rank}(H^{(k)})$ for each cell via SVD. Condition (G1) is supported if at least $d_{\mathrm{eff}}+1$ cells achieve rank $|\mathcal{F}|-1$, where $d_{\mathrm{eff}}$ is the effective dimension of the gate feature space. Among the cells satisfying (G1), check that their representative feature values span $\mathbb{R}^{d_{\mathrm{eff}}}$ affinely; (G2) is supported if this affine rank condition holds.
@@
Because the gate features $z(A)$ are $12$-dimensional summary statistics, many distinct menus share the same (or very close) feature value. We group menus into cells $\mathcal{C}_{k}$ using exact equality of $z(A)$ where possible, supplemented by $k$-means clustering as a finite-sample heuristic. We retain cells with at least $|\mathcal{F}|-1$ menus. For each cell, we form the matrix $\hat{H}^{(k)}$ by stacking the row vectors $\hat{h}(A)^{\top}=\big(\kappa_{f}^{L}(A)-\hat{r}(A)\kappa_{f}^{R}(A)\big)_{f\in\mathcal{F}}$ over its menus (trimming $\hat{p}(A)$ to $\varepsilon,1-\varepsilon]$ for numerical stability), and compute its numerical rank via SVD with a tolerance-based criterion. Assumption (G1) is supported if at least $d_{\mathrm{eff}}{+}1$ cells achieve $\mathrm{rank}(\hat{H}^{(k)})=|\mathcal{F}|-1$. To verify (G2), we check that the centroids of cells satisfying (G1) are affinely independent.
@@
D1–D2 document that the model generates informative variation in left/right decisive sets. D3 provides empirical support for the sufficient conditions of Theorem 1 under the baseline gate specification with $z(A)$. We report the empirical results of these diagnostics on the Choices13k dataset in Section 9.1.
@@ 4 Rule library
The outcomes $x,y$ enter in their raw units (before the outcome rescaling applied to neural network features in Section 6.3). The constant +1+1 in the denominator is a fixed regularization choice (not a tunable parameter) that prevents division by zero and pins down the scale in the raw-outcome units of the dataset. Because the contrast is not homogeneous of degree zero, the effective rule recommendations depend on the outcome scale of the data. The rules remain parameter-free in the sense that no parameter is estimated from data, but they are not invariant to affine transformations of outcomes; applying the same rules to payoffs denominated in cents rather than dollars could change which rules are active on some menus. This scale dependence is a limitation of the current formulation. For some rules we use a small set of extreme pairings. For $A=\{L^{1},L^{2}\}$ define
@@
which contains at most four outcome pairs.
@@ 4.1 Rule-level Diagnostics
$w_{f}:=\frac{1}{N}\sum_{i=1}^{N}\tilde{q}_{if},\qquad\sum_{f\in\mathcal{F}}w_{f}=1.$ (14)

### 19 Mar 2026 v2

@@ 2.1 Setup
A menu (choice set) is a binary set $A=\{L^{1},L^{2}\}$ of lotteries, each lottery being a finite-support distribution over monetary outcomes. We label $L^{1}$ the left choice and $L^{2}$ the right choice. Let $\mathcal{A}_{\mathrm{full}}$ denote the full set of menus in the dataset. Each $A\in\mathcal{A}_{\mathrm{full}}$ is observed $n(A)\geq 1$ times. Let $c_{m}(A)\in\{L^{1},L^{2}\}$ be the realized choice in repetition $m\in\{1,\dots,n(A)\}$. The menu-level left-choice frequency is
$\hat{p}(A)\ :=\ \frac{1}{n(A)}\sum_{m=1}^{n(A)}\mathbf{1}\{c_{m}(A)=L^{1}\}\in[0,1].$
@@
The primitives of our model are lotteries, and all subsequent rules - including regret- and disappointment-type rules - are defined as functions of the lottery representation (or of a canonical state space constructed from the marginal distributions via the product measure over outcome pairs, as described in Section 4). Hence the lottery representation is without loss of generality for the properties we study: “without loss” is relative to the paper’s rule definitions, which take lottery representations as inputs, not joint state-act distributions. [^1] For the regret- and disappointment-type rules, we construct a canonical state space from the product measure over outcome pairs of the two lotteries (Section 4). This coupling is a specific construction that does not invoke or require data on correlated state-act outcomes. We will use the state-space view only when defining several rules, and otherwise work with the simpler lottery notation introduced above.
@@ 3 Identification via Rule Switching
This section clarifies when the gate parameters $\theta=\{(\alpha_{f},\beta_{f})\}_{f\in\mathcal{F}}$ are pinned down by menu-level choice frequencies. The identification argument has two layers. The first layer identifies, at a fixed feature value, the vector of relative rule weights. The second layer uses variation across feature values to recover the affine gate parameters. The key identifying force throughout is rule switching: because different rules recommend different sides in different menus, the decisive-side pattern varies across the menu support, and this variation pins down the latent rule weights.
From CA-RRM to a collapsed binary logit index.
@@
Let
$\mathcal{A}_{2}:=\{A\in\mathcal{A}_{\mathrm{full}}:\ \mathcal{I}_{L}(A)\neq\emptyset,\ \mathcal{I}_{R}(A)\neq\emptyset\}$
denote the set of two-sided menus. For $A\in\mathcal{A}_{2}$, define the population log-odds
@@
Let $\vartheta\in\mathbb{R}^{K}$ collect the remaining free parameters, with
$K=(|\mathcal{F}|-1)\,(1+\dim(\psi)).$
@@
and write the odds ratio
$r(A):=\frac{p(A)}{1-p(A)}.$
For any $x\in\mathbb{R}^{d}$, define the positive weight vector
$\omega_{f}(x):=\exp(\alpha_{f}+\beta_{f}^{\top}x).$
Cross-multiplying the odds form of the model yields, for each menu $A$ with $\psi(A)=x$,
@@
i.e. $h(A)^{\top}\omega(x)=0$, where
$h_{f}(A):=\kappa_{f}^{L}(A)-r(A)\kappa_{f}^{R}(A).$
Identification in two layers.
Equation (11) reveals the first identifying object: at a fixed feature value $x$, every menu with $\psi(A)=x$ shares the same latent weight vector $\omega(x)$, but may display a different decisive-side pattern. These distinct decisive-side patterns generate linear restrictions on the same unknown vector $\omega(x)$. Once $\omega(x)$ is identified at sufficiently many feature values, the affine structure of the gate recovers $(\alpha_{f},\beta_{f})$ rule by rule.
Theorem 1 (Fixed-feature identification of rule weights).
Fix $x\in\mathbb{R}^{d}$ and define
$\mathcal{A}_{2}(x):=\{A\in\mathcal{A}_{2}:\psi(A)=x\}.$
For each $A\in\mathcal{A}_{2}(x)$, let
$h_{f}(A):=\kappa_{f}^{L}(A)-r(A)\kappa_{f}^{R}(A),\qquad r(A):=\frac{p(A)}{1-p(A)}.$
Define the cellwise identified set
$\Omega(x):=\Big\{\omega\in\mathbb{R}_{++}^{|\mathcal{F}|}:h(A)^{\top}\omega=0\ \ \forall A\in\mathcal{A}_{2}(x)\Big\}.$
Assume $\Omega(x)\neq\varnothing$ (this is guaranteed under correct specification: the true weights $\omega_{f}(x)=\exp(\alpha_{f}+\beta_{f}^{\top}x)$ are strictly positive by construction, so $\omega(x)\in\Omega(x)$). Then the following are equivalent:
  1. (i)
    $\Omega(x)$ is a single positive ray, i.e. there exists $\omega^{\star}(x)\in\mathbb{R}_{++}^{|\mathcal{F}|}$ such that
    $\Omega(x)=\{c\,\omega^{\star}(x):c>0\};$
  2. (ii)
    the row span of $\{h(A):A\in\mathcal{A}_{2}(x)\}$ has dimension $|\mathcal{F}|-1$.
Equivalently, if $H(x)$ is any stacked matrix whose rows span the same subspace as $\{h(A):A\in\mathcal{A}_{2}(x)\}$, then
$\Omega(x)\text{ is a single positive ray}\quad\Longleftrightarrow\quad\operatorname{rank}H(x)=|\mathcal{F}|-1.$
Under the normalization $\omega_{f_{0}}(x)=1$, this is equivalent to point identification of $\omega(x)$.
Proof sketch.
The model implies $h(A)^{\top}\omega(x)=0$ for every $A\in\mathcal{A}_{2}(x)$, so $\Omega(x)$ is the set of strictly positive vectors in the null space of the stacked matrix $H(x)$. If $\operatorname{rank}H(x)=|\mathcal{F}|-1$, the null space is one-dimensional and $\Omega(x)$ is a single positive ray; the converse follows because, if the null space has dimension at least two and already contains one strictly positive vector (namely the true $\omega(x)$), then a sufficiently small perturbation along an independent null-space direction yields a second, non-proportional positive vector that still lies in the positive orthant. See Appendix B for the full proof. ∎
The key identifying force in Theorem 1 is rule switching at fixed features: menus sharing the same $\psi(A)=x$ share the same latent weight vector $\omega(x)$, but may differ in which rules recommend which option. Each such menu contributes a linear restriction $h(A)^{\top}\omega(x)=0$, and identification occurs when these restrictions span a codimension-one subspace of $\mathbb{R}^{|\mathcal{F}|}$.
From fixed-feature weights to global identification of the affine gate.
Theorem 1 identifies the first-stage object $\omega(x)$ feature value by feature value. The next result shows that once these weight vectors are recovered at sufficiently many affinely independent feature values, the affine gate parameters themselves are globally identified.
Corollary 2 (Global identification from fixed-feature weights).
Consider the CA-RRM with affine indices
$s_{f}(A)=\alpha_{f}+\beta_{f}^{\top}\psi(A),\qquad f\in\mathcal{F},$
and impose the normalization
$(\alpha_{f_{0}},\beta_{f_{0}})=(0,0)$
for some baseline rule $f_{0}\in\mathcal{F}$. Let $d=\dim(\psi)$. Suppose there exist $(d+1)$ feature values $x^{(0)},\ldots,x^{(d)}$ such that:
@@
    Cellwise identification: For each $k\in\{0,\ldots,d\}$, there exist menus $A_{k1},\ldots,A_{kM_{k}}$ with $\psi(A_{km})=x^{(k)}$ for all $m$, such that the associated matrix
@@
    Affine richness of feature values: The matrix
    $X:=\begin{pmatrix}1&(x^{(0)})^{\top}\\ \vdots&\vdots\\ 1&(x^{(d)})^{\top}\end{pmatrix}$
    has full rank $d+1$.
@@
Proof sketch.
By Theorem 1, (G1) identifies each $\omega(x^{(k)})$ up to scale; the normalization $\omega_{f_{0}}\equiv 1$ pins the scale. For each rule $f$, stacking $\log\omega_{f}(x^{(k)})=\alpha_{f}+\beta_{f}^{\top}x^{(k)}$ over $k=0,\ldots,d$ gives a linear system with coefficient matrix $X$, which is invertible by (G2). See Appendix C for the detailed version. ∎
The corollary separates global identification into two conceptually distinct requirements. Condition (G1) is a within-feature condition: it uses rule switching across menus that share the same feature value to identify the relative rule weights at that feature value. Condition (G2) is an across-feature condition: it uses affine variation in the feature values themselves to map the identified weight vectors back into the gate coefficients. Thus, the identification argument has a clear two-step structure: first recover $\omega(x)$ locally in feature space, then recover the affine law of motion of those weights globally.
Effective dimension and baseline specification.
Corollary 2 is stated for a generic feature map $\psi$. In our baseline specification (Section 6.6), the gate operates on a $12$-dimensional vector of interpretable summary statistics $z(A)\in\mathbb{R}^{12}$, so the corollary applies with $\psi=z$. When this feature vector has structural redundancies (e.g., exact linear dependencies among coordinates), the corollary applies with $d$ replaced by the effective dimension
$d_{\mathrm{eff}}:=\operatorname{rank}([\mathbf{1},z])-1,$
and identification holds for the gate parameters projected onto the effective feature subspace (i.e., up to the null space of $\mathbf{1},z]$). In our application, $d_{\mathrm{eff}}=11$ (the expected-value gap is a linear combination of the two level expected values). The neural benchmark models use a separate, higher-dimensional raw menu encoding $\psi(A)\in\mathbb{R}^{4M}$ as their input (Section 6.3); the identification conditions and diagnostics below do not target that encoding.
@@
Theorem 1 and Corollary 2 reduce global identification to two verifiable ingredients on the observed support: cellwise rank sufficiency and affine richness of the identifying feature values. We report three diagnostics.
@@
Report the fraction of menus that are two-sided:
$\widehat{\Pr}(A\in\mathcal{A}_{2}).$
This ensures that the log-odds representation (9) is empirically meaningful.
@@
The diagnostic is implemented on the interpretable gate features $z(A)\in\mathbb{R}^{12}$ used in the baseline specification (Section 6.6). To check (G1), group menus into cells that share the same (or nearby) value of $z(A)$, and compute $\operatorname{rank}(H^{(k)})$ for each cell via singular value decomposition (SVD). Condition (G1) is supported if at least $d_{\mathrm{eff}}+1$ cells achieve rank $|\mathcal{F}|-1$, where $d_{\mathrm{eff}}$ is the effective dimension of the gate feature space. Among the cells satisfying (G1), check that their representative feature values span $\mathbb{R}^{d_{\mathrm{eff}}}$ affinely; (G2) is supported if this affine rank condition holds.
@@
Because the gate features $z(A)$ are $12$-dimensional summary statistics, many distinct menus share the same (or very close) feature value. We group menus into cells $\mathcal{C}_{k}$ using exact equality of $z(A)$ where possible, supplemented by $k$-means clustering as a finite-sample heuristic. We retain cells with at least $|\mathcal{F}|-1$ menus. For each cell, we form the matrix
$\hat{H}^{(k)}$
by stacking the row vectors
$\hat{h}(A)^{\top}=\big(\kappa_{f}^{L}(A)-\hat{r}(A)\kappa_{f}^{R}(A)\big)_{f\in\mathcal{F}}$
over its menus (trimming $\hat{p}(A)$ to $\varepsilon,1-\varepsilon]$ for numerical stability), and compute its numerical rank via SVD. [^3] With estimated $\hat{r}(A)$, the matrix $\hat{H}^{(k)}$ generically has full algebraic column rank $|\mathcal{F}|$ once the cell contains enough menus. We therefore check approximate rank deficiency by retaining singular values above a tolerance $\tau=10^{-8}\times\sigma_{1}(\hat{H}^{(k)})$, where $\sigma_{1}$ is the largest singular value. (G1) is supported when the gap between $\sigma_{|\mathcal{F}|-1}$ and $\sigma_{|\mathcal{F}|}$ is large relative to $\tau$, so that the rank-$(|\mathcal{F}|-1)$ structure is robust to perturbations from sampling noise in $\hat{r}(A)$. Assumption (G1) is supported if at least $d_{\mathrm{eff}}+1$ cells achieve
$\operatorname{rank}(\hat{H}^{(k)})=|\mathcal{F}|-1.$
To verify (G2), we check that the centroids of cells satisfying (G1) are affinely independent.
@@
D1–D2 document that the model generates informative variation in left/right decisive sets. D3 provides empirical support for the sufficient conditions of Corollary 2 under the baseline gate specification with $z(A)$. We report the empirical results of these diagnostics on the Choices13k dataset in Section 7.1.
@@ 4 Rule library
The outcomes $x,y$ enter in their raw units (before the outcome rescaling applied to neural network features in Section 6.3). The constant +1+1 in the denominator is a fixed regularization choice (not a tunable parameter). Because the contrast is not homogeneous of degree zero, the effective rule recommendations depend on the outcome scale of the data. The rules remain parameter-free in the sense that no parameter is estimated from data, but they are not invariant to affine transformations of outcomes; applying the same rules to payoffs denominated in cents rather than dollars could change which rules are active on some menus. This scale dependence is a limitation of the current formulation. For some rules we use a small set of extreme pairings. For $A=\{L^{1},L^{2}\}$ define
@@
which contains at most four outcome pairs. [^4] When $x_{\min}(L^{i})=x_{\max}(L^{i})$ for some lottery, $\mathcal{P}(A)$ may contain fewer than four distinct pairs. The notation $\{(a_{k},b_{k})\}_{k=1}^{4}$ should be understood as an indexed enumeration of the $2\times 2$ grid (allowing repeats); the salience-score and ordering operations below remain well-defined in all cases.
@@ 4.1 Rule-level Diagnostics
$w_{f}:=\frac{1}{T}\sum_{t=1}^{T}\tilde{q}_{tf},\qquad\sum_{f\in\mathcal{F}}w_{f}=1.$ (14)